2. Solve 2x + 3y = 11 and 2x – 4y = – 24 and hence find the value of ‘m’ for which y = mx + 3.

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
2x + 3y = 11…………………………..(I) 2x – 4y = -24………………………… (II) From equation (II), we get x = (11-3y)/2 ………………….(III) Substituting the value of x in equation (II), we get 2(11-3y)/2 – 4y = 24 11 – 3y – 4y = -24 -7y = -35 y = 5……………………………………..(IV) Putting the value of y in equation (III), we get x = (11-3×5)/2 = -4/2 = -2 Hence, x = -2, y = 5 Also, y = mx + 3 5 = -2m +3 -2m = 2 m = -1